1.two numbers,which differ by 3,have a product off 88.find them.
2.A man walks a certain ditance due north and then the same distance plus a further 7km de east.if the final distance from the stating point is 17km,find the distances he walks north to east.
3.A farmer makes a profit of x cents on each of the(x+5)eggs her hen lays.if her total profit was 84 cents,find the number of eggs the hen lays.
4.A boy buys x eggs at (x+8) cents each and (x-2) rashers of bacon at (x-3) cents each. if the total bills is $1.75, how many egs does he buy?

Answer by fractalier(6550)

I see no one has answered this, because you need to submit them one at a time. I will do the first one...please resubmit the others.
Call the numbers x and y. Then we have
x = y - 3
We also have
xy = 88
Now we substitute and get
(y - 3)y = 88
y^2 - 3y - 88 = 0
(y - 11)(y + 8) = 0
y = 11 and x = 8
or
y = -8 and x = -11